OpenStudy (katecc379):

Simonne has been following the stock ratings of Carter's Pork Bellies. She has noticed that the pattern can be modeled by a cosine function. During her observations, she has found the stock to have a maximum of x = 6, minimum of x = −2, and period of π. Which of the following cosine functions could model Simonne's stock rating?

OpenStudy (katecc379):

@UnkleRhaukus

OpenStudy (agent0smith):

\[\large x = A \cos k t + c\] First find the midline, which will be c - it's halfway between the max and min. Then the amplitude, A - distance between midline and max (or min) And to find k, use: period = \(\Large \frac{ 2\pi }{ k }\)